# Machine Vision-Based Assessment of Fall Color Changes in Apple Leaves and its Relationship with Nitrogen Concentration

### v3

@@ 1 Introduction
Water and nutrient applications in commercial tree fruit orchards are managed at a block or field level, providing treatment relatively uniformly to hundreds or even thousands of trees without considering the variability among individual trees. This homogeneous management of nutrients in tree fruit crops often leads to sub- and/or super-optimal inputs, with negative impacts on tree growth, fruit yield, and fruit quality (Klein et al., 1989; Neilsen et al., 2009). Managing these inputs at the individual tree level (i.e. addressing the unique needs of individual trees) is a promising concept for improving tree vigor, fruit yield and quality. However, managing orchards (Figure 1) at the individual tree level is challenging, particularly in terms of assessing the needs of individual trees and making precise management decisions. Each tree within an orchard is unique with its specific needs for water, nutrients, and other inputs (Aggelopoulou et al., 2010; James A. Taylor et al., 2007). For eg., Aggelopoulou et al. (2010) reported spatial variability with a coefficient of variation (CV) of up to 82% for bloom and 40% for yield within the same block. The variability among trees can be caused by multiple factors including spatial variation of soil properties like texture, depth, slope, or microbial activities, variation in microclimates within orchards, and differences in tree training and pruning (Aggelopoulou et al., 2010, 2011; Umali et al., 2012; Cho, 2010; Parkin, 1993). Effective nutrient management at the individual tree level could be possible if these multiple factors can be accurately assessed for each individual tree.
@@
Machine vision techniques have been extensively used to evaluate various canopy-level features and detect canopy stresses in fruit crops (Wang et al., 2023; Paudel et al., 2022; Wachs et al., 2010; Wang et al., 2018; Brown et al., 2024). More specifically, many recent studies have reported machine vision-based tools for estimating canopy density(Mahmud et al., 2021; Paudel et al., 2023), tree trunk cross-sectional area (Wang et al., 2023), and canopy color (Naschitz et al., 2014) in fruit trees.
Leaf color is one of the key canopy characteristics that is correlated with nitrogen content(Ali et al., 2012; Treder et al., 2016; Ye et al., 2020). This correlation is due to the presence of four rings of N in the chlorophyll, the pigment in leaves responsible for light harvesting(Marschner, 2011). Chlorophyll reflects light primarily in the green spectrum, which is why plant tissue containing chlorophyll appears green. In deciduous trees like apples, a significant color change occurs from green to yellow as the growing season progresses towards fall, as the chlorophyll breaks down and mobile elements like N move back into the perennial structures of tree. This relocation of nitrogen from leaves to woody regions for storage will later support next seasons’ growth(Murneek and Logan, 1932; Spencer, 1973; Neilsen and Neilsen, 2002).A greener canopy with vigorous growth is often correlated with a tree having high nitrogen, while a yellower and less vigorous canopy is often associated with trees with low nitrogen (Raese et al., 2007; Lee, 2002). In addition, trees with high nitrogen tend to maintain their green color longer into the season providing a potential indication of tree nitrogen status during this critical transition.
Various RGB, multispectral, and hyperspectral imaging techniques have been evaluated to quantify the color and estimate nitrogen or chlorophyll content in apples (Campbell et al., 1990; Neilsen et al., 1995; Perry and Davenport, 2007). Most of these studies use devices like chlorophyllmeter (SPAD) or satellite images. While handheld devices like SPAD meters can be utilized in the field, they pose challenges for scaling assessments across entire orchards due to the need for data collection from individual leaves. On the other hand, satellite imagery offers broader insights but often lacks the resolution required to pinpoint individual trees, necessitating further computations for accurate matching. Our study overcomes these limitations by focusing on individual tree-level assessments, providing a scalable and efficient approach for evaluating apple tree health and productivity in orchards.
@@ 2 Materials and Methods
RGB-D images were collected over several weeks in fall of 2021 and 2023. These images were collected in a Jazz commercial orchard (Yakima Valley Orchard) near Prosser, WA, USA (Figure 3). The trees were trained to a vertical structure using a tall spindle canopy architecture (Figure 1). The point clouds of the tree canopies were then segmented into yellow and green foliage; a quantitative metric, termed the yellowness index, was created to estimate the color of the trees. The yellowness indices of the trees during different weeks were then correlated with the leaf nitrogen content for each year. The overall data collection and processing method used in this study is shown in Figure 2.
@@ 2.1 Data Collection
The datasets (color and 3D images) for this study were collected over five weeks in fall, 2021 (from October 13 to November 14) and fall, 2023 (from October 20 to November 17), every week. The first image datasets were determined by when the leaves began to change color from green to yellow (Figure 4), The dataset was collected from a total of 200 trees across 17 rows of the test orchard, with every other 3rd tree in each row taken as a sample. The data were collected using a commercial stereovision-based RGB-D camera (Zed2i, Stereolabs, Paris, France) mounted on a utility vehicle (Figure 3b). The camera was mounted  2 meters above the ground to ensure that most of the canopy (Figure 4, part of tree above  0.3m from ground) was visible. An SVO file (Stereolabs file format) with all the metadata including the camera parameters, images, and sensor data was recorded for each row of trees. Images (pixel resolution of 2208 x 1242) and point clouds (maximum depth of 10m) for each tree were later extracted based on timestamp matching between frames and user input. During processing, the RGB and co-registered point cloud were identified for each test tree.
@@ 2.2.1 Foreground Tree Segmentation
Figure 5: Schematic for point cloud segmentation and downsampling process used in the study
@@ 2.2.2 Color Segmentation and Clustering
The point cloud obtained from step 2.2.1 consisted of points from the foreground tree, and could be broadly categorized into yellow leaves, green leaves, and the trunk. To quantify the color changes in trees’ over the duration of study, HSV (Figure 6a) and CIE-Lab* (Figure 6b and 6c) color spaces were used as they are less affected by the variation in intensities. In the HSV space, hue remains unaffected by varying light intensity, while in the CIE-Lab* space, the a* and b* values correspond respectively to red-green and blue-yellow colors. RGB images captured by the RGB-D sensor were converted to both the CIE-La*b* (Lab) space, using a standard D65 illuminant, and the HSV color space. These conversions were performed using the scikit-image library (Van der Walt et al., 2014).
The transition of the trees from green to yellow color was noticeable in both HSV and La*b* space (Figures 6a and 6b,c). A probability density plot for the hue space across the weeks illustrated a decline from approximately 110 degrees in week 1 to about 50 degrees in week 6 (hue expressed as degrees between 0 and 360, as shown in Figure 6a), signifying the shift from green to yellow. A similar trend was visible in the probability density plots for a* and b* in the La*b* space. Both a* and b* showed a slight increase as the weeks progressed, corresponding to a color change from green to yellow.
@@
The transition in tree canopy color was apparent qualitatively, but a robust quantitative approach was needed to precisely quantify them. To achieve this, individual points in the point cloud were categorized into distinct groups of yellow leaves, green leaves, or the trunk. Qualitative analyses using K-means clustering over multiple samples in both the La*b* and HSV color spaces showed that the La*b* space performed better in distinguishing yellow and green pixels as separate classes. Consequently, the subsequent analysis was conducted using La*b* color space. To cluster the point cloud, a modified k-means algorithm (Algorithm 1) was applied, initially grouping the points into 20 clusters. A threshold for the cluster center was set in both the a* and b* space, to merge these clusters into three distinct classes: $Yellow$, $Green$, and $Trunk$. The $Yellow$ cluster encompassed foliage that had transitioned to yellow, while the $Green$ cluster comprised foliage that retained its green color. The $Trunk$ cluster included points from the trunk, branches, background soil, and some brown leaves. A detailed breakdown of the clustering process is outlined in Algorithm 1.
@@ 20                     $c_{t}=c_{t}\cup c_{i}$
An alternative approach utilized a gradient boost classifier, an ensemble learning technique that combines multiple weak learners (typically decision trees) to create a stronger, more robust model (Bartlett et al., 1998).. It sequentially builds new models, focusing on correcting the errors made by the previous ones, ultimately producing a more accurate and powerful predictive model (Natekin and Knoll, 2013). In this study, the gradient boost classifier model was created using scikit-learn (Van der Walt et al., 2014). The dataset was divided into 80% training and 20% testing sub-sets. The model required tuning of different hyperparameters. A strategic approach to varying model hyperparameters was employed, with one hyperparameter held constant while varying the others to analyze their individual effects on model accuracy. The hyperparameter learning rate was varied between 0.1-1, max depth was varied between 1-5, and n estimator was varied between 100-1000 and their performance on training and test dataset was analyzed. The set of hyperparameters that achieved the highest accuracy on the test data was then selected.
@@ 2.3 Yellowness Index and Leaf Nitrogen Concentration
To investigate if the yellowness index had any relationship with the foliar nitrogen concentration on the trees, the yellowness index values were regressed against the leaf nitrogen concentration. The leaf nitrogen concentration was obtained using the standard Kjeldahl’s method(Kirk, 1950) by collecting 50 fully grown mid-shoot leaves from each tree, at the end of the growing season (Late July in 2021 and early August in 2023). The leaf nitrogen concentration represents the average leaf nitrogen concentration of each tree and is associated with canopy growth and canopy color. The leaf nitrogen level of 2-2.4% was regarded as an adequate range for nitrogen concentration based on referenced values for apples(Cheng and Schupp, 2004; Cheng, 2010; Sallato, 2017; UniversityofArizona, ).
@@ 3.1 Tree Segmentation and Clustering
Figure 13: The images (top row) and segmented point cloud (bottom row) of a test tree during weeks 1,3, and 5. The three colors in the point cloud represent three groups; green-green leaves, yellow-yellow leaves, and brown-trunk, branches and some leaves that have turned brown.
@@ 3.2 Yellowness Index and Leaf Nitrogen Concentration
The correlation between leaf nitrogen concentration and yellowness index was determined for each week in 2021 and 2023 (Figure 15). The trees were categorized into five groups based on their leaf nitrogen concentrations as outlined in Section 2.3. In general, tree nitrogen levels were poorly correlated with yellowness index in both years (r values ranging from -0.28 to -0.43 in 2021 and 0.16 and -0.64 in 2023). We also observed an extended retention of greenness in trees that had high N content compared to those with lower N levels. Previous studies have shown strong correlations between color indices and leaf nitrogen concentration in various plant species. For example, Lee et al. (2019) reported a correlation coefficient of 0.85-0.91 between the SPAD readings and leaf Nitrogen concentration in apples, while El-Azazy (2018) found an $R^{2}$ of 0.81 for citrus trees. Haider et al. (2021) observed an even higher $R^{2}$ of 0.91 across multiple crops, though their system required crop-specific calibration. Our study’s lower correlation might be related to two factors: variable lighting conditions in natural settings, and a limited range of samples in the lower nitrogen concentration region.
@@
The results from this study demonstrated that a machine vision system could be used to identify differences in the yellowing patterns of trees. The outcomes and methodologies of this study offer a means to quantify tree color that could be used to determine tree nitrogen levels, which are crucial factors in guiding fertilizer application to individual trees. However, the system relies on a machine vision system which is susceptible to variations in external lighting conditions. Nevertheless, recent advancements in machine vision sensors have addressed low-light conditions using innovative techniques such as high dynamic range (Zhang et al., 2020, 2021) and Adversarial Networks(Chen et al., 2018). In addition, deep learning methods have notably enhanced imaging system resilience to diverse lighting environments(Li et al., 2021a, b). The algorithm employed in this study relies on labeled data for initially training the model to differentiate yellow and green points on trees, necessitating manual inputs. It’s important to note that this algorithm might require further calibration between different years and across tree varieties for optimal performance.

### v4

@@ 1 Introduction
Water and nutrient applications in commercial tree fruit orchards are managed at a block or field level, providing treatment relatively uniformly to hundreds or even thousands of trees without considering the variability among individual trees. This homogeneous management of nutrients in tree fruit crops often leads to sub- and/or super-optimal inputs, with negative impacts on tree growth, fruit yield, and fruit quality (Klein et al., 1989; Neilsen et al., 2009). Managing these inputs at the individual tree level (i.e. addressing the unique needs of individual trees) is a promising concept for improving tree vigor, fruit yield and quality. However, managing orchards at the individual tree level is challenging, particularly in terms of assessing the needs of individual trees and making precise management decisions. Each tree within an orchard is unique with its specific needs for water, nutrients, and other inputs (Aggelopoulou et al., 2010; James A. Taylor et al., 2007). For eg., Aggelopoulou et al. (2010) reported spatial variability with a coefficient of variation (CV) of up to 82% for bloom and 40% for yield within the same block. The variability among trees can be caused by multiple factors including spatial variation of soil properties like texture, depth, slope, or microbial activities, variation in microclimates within orchards, and differences in tree training and pruning (Aggelopoulou et al., 2010). Effective nutrient management at the individual tree level could be possible if these multiple factors can be accurately assessed for each individual tree.
@@
Machine vision techniques have been extensively used to evaluate various canopy-level features and detect canopy stresses in fruit crops (Wang et al., 2023; Paudel et al., 2022; Wang et al., 2018; Brown et al., 2024). More specifically, many recent studies have reported machine vision-based tools for estimating canopy density(Mahmud et al., 2021; Paudel et al., 2023), tree trunk cross-sectional area (Wang et al., 2023), and canopy color (Naschitz et al., 2014) in fruit trees.
Leaf color is one of the key canopy characteristics that is correlated with nitrogen content(Ali et al., 2012; Ye et al., 2020). This correlation is due to the presence of four rings of N in the chlorophyll, the pigment in leaves responsible for light harvesting(Marschner, 2011). Chlorophyll reflects light primarily in the green spectrum, which is why plant tissue containing chlorophyll appears green. In deciduous trees like apples, a significant color change occurs from green to yellow as the growing season progresses towards fall, as the chlorophyll breaks down and mobile elements like N move back into the perennial structures of tree. This relocation of nitrogen from leaves to woody regions for storage will later support next seasons’ growth(Murneek and Logan, 1932; Neilsen and Neilsen, 2002).A greener canopy with vigorous growth is often correlated with a tree having high nitrogen, while a yellower and less vigorous canopy is often associated with trees with low nitrogen (Raese et al., 2007). In addition, trees with high nitrogen tend to maintain their green color longer into the season providing a potential indication of tree nitrogen status during this critical transition.
Various RGB, multispectral, and hyperspectral imaging techniques have been evaluated to quantify the color and estimate nitrogen or chlorophyll content in apples (Neilsen et al., 1995; Perry and Davenport, 2007). Most of these studies use devices like chlorophyllmeter (SPAD) or satellite images. While handheld devices like SPAD meters can be utilized in the field, they pose challenges for scaling assessments across entire orchards due to the need for data collection from individual leaves. On the other hand, satellite imagery offers broader insights but often lacks the resolution required to pinpoint individual trees, necessitating further computations for accurate matching. Our study overcomes these limitations by focusing on individual tree-level assessments, providing a scalable and efficient approach for evaluating apple tree health and productivity in orchards.
@@ 2 Materials and Methods
RGB-D images were collected over several weeks in fall of 2021 and 2023. These images were collected in a Jazz commercial orchard (Yakima Valley Orchard) near Prosser, WA, USA.Data was collected from a commercial apple orchard located in Prosser, WA (Figure 1), where jazz apple trees were grafted onto M9 rootstock. The trees were planted in 2008 and trained using a tall spindle, vertical fruiting wall system. They were spaced closely at 1.4 meters (4.5 feet) apart, with 2.6 meters (8.5 feet) between adjacent rows. The irrigation system included a combination of drip lines and both overhead and under-canopy irrigation, laid out uniformly throughout. The soil composition consists of silt loam over a layer of calcium carbonate, approximately 1-2 inches thick, which sits atop basalt. The point clouds of the tree canopies were then segmented into yellow and green foliage; a quantitative metric, termed the yellowness index, was created to estimate the color of the trees. The yellowness indices of the trees during different weeks were then correlated with the leaf nitrogen content for each year. The overall data collection and processing method used in this study is shown in Figure 2.
@@ 2.1 Data Collection
The datasets (color and 3D images) for this study were collected over five weeks in fall, 2021 (from October 13 to November 14) and fall, 2023 (from October 20 to November 17), every week. The time to start the collection of data was determined by when the leaves began to change color from green to yellow (Figure 4) upon consultation with the orchard manager. The timing of this is dependent on a variety of factor including temperature throughout the growing season, the location of the orchard, and other management practices in the orchard. The dataset was collected from a total of 200 trees across 17 rows of the test orchard, with every other 3rd tree in each row taken as a sample. The data were collected using a commercial stereovision-based RGB-D camera (Zed2i, Stereolabs, Paris, France) mounted on a utility vehicle (Figure 3b). The camera was mounted  2 meters above the ground to ensure that most of the canopy (Figure 4, part of tree above  0.3m from ground) was visible. An SVO file (Stereolabs file format) with all the metadata including the camera parameters, images, and sensor data was recorded for each row of trees. Images (pixel resolution of 2208 x 1242) and point clouds (maximum depth of 10m) for each tree were later extracted based on timestamp matching between frames and user input. During processing, the RGB and co-registered point cloud were identified for each test tree.
@@ 2.2.2 Color Segmentation and Clustering
The point cloud obtained from step 2.2.1 consisted of points from the foreground tree, and could be broadly categorized into yellow leaves, green leaves, and the trunk. To quantify the color changes in trees’ over the duration of study, HSV (Figure 5a) and CIE-Lab* (Figure 5b and 5c) color spaces were used as they are less affected by the variation in intensities. In the HSV space, hue remains unaffected by varying light intensity, while in the CIE-Lab* space, the a* and b* values correspond respectively to red-green and blue-yellow colors. RGB images captured by the RGB-D sensor were converted to both the CIE-La*b* (Lab) space, using a standard D65 illuminant, and the HSV color space. These conversions were performed using the scikit-image library (Van der Walt et al., 2014).
The transition of the trees from green to yellow color was noticeable in both HSV and La*b* space (Figures 5a and 5b,c). A probability density plot for the hue space across the weeks illustrated a decline from approximately 110 degrees in week 1 to about 50 degrees in week 6 (hue expressed as degrees between 0 and 360, as shown in Figure 5a), signifying the shift from green to yellow. A similar trend was visible in the probability density plots for a* and b* in the La*b* space. Both a* and b* showed a slight increase as the weeks progressed, corresponding to a color change from green to yellow.
@@
The transition in tree canopy color was apparent qualitatively, but a robust quantitative approach was needed to precisely quantify them. To achieve this, individual points in the point cloud were categorized into distinct groups of yellow leaves, green leaves, or the trunk. Two different approaches were used: a) An unsupervised K-means based method and b) A supervised Gradient Boosting Method.
K-Means Based Segmentation: Qualitative analyses using K-means clustering over multiple samples in both the La*b* and HSV color spaces showed that the La*b* space performed better in distinguishing yellow and green pixels as separate classes. Consequently, the subsequent analysis was conducted using La*b* color space. To cluster the point cloud, a modified k-means algorithm (Algorithm 1) was applied, initially grouping the points into 20 clusters. A threshold for the cluster center was set in both the a* and b* space, to merge these clusters into three distinct classes: $Yellow$, $Green$, and $Trunk$. The $Yellow$ cluster encompassed foliage that had transitioned to yellow, while the $Green$ cluster comprised foliage that retained its green color. The $Trunk$ cluster included points from the trunk, branches, background soil, and some brown leaves. A detailed breakdown of the clustering process is outlined in Algorithm 1.
@@ 20                     $c_{t}=c_{t}\cup c_{i}$
Gradient Boost Classifier Based Segmentation: An alternative approach utilized a gradient boost classifier, an ensemble learning technique that combines multiple weak learners (typically decision trees) to create a stronger, more robust model (Bartlett et al., 1998). It sequentially builds new models, focusing on correcting the errors made by the previous ones, ultimately producing a more accurate and powerful predictive model. In this study, the gradient boost classifier model was created using scikit-learn (Van der Walt et al., 2014). The dataset was divided into 80% training and 20% testing sub-sets. The model required tuning of different hyperparameters. A strategic approach to varying model hyperparameters was employed, with one hyperparameter held constant while varying the others to analyze their individual effects on model accuracy. The hyperparameter learning rate was varied between 0.1-1, max depth was varied between 1-5, and n estimator was varied between 100-1000 and their performance on training and test dataset was analyzed. The set of hyperparameters that achieved the highest accuracy on the test data was then selected.
@@ 2.3 Yellowness Index and Leaf Nitrogen Concentration
To investigate if the yellowness index had any relationship with the foliar nitrogen concentration on the trees, the yellowness index values were regressed against the leaf nitrogen concentration. The leaf nitrogen concentration was obtained using the standard Kjeldahl’s method(Kirk, 1950) by collecting 50 fully grown mid-shoot leaves from each tree, at the end of the growing season (Late July in 2021 and early August in 2023). The leaf nitrogen concentration represents the average leaf nitrogen concentration of each tree and is associated with canopy growth and canopy color. The leaf nitrogen level of 2-2.4% was regarded as an adequate range for nitrogen concentration based on referenced values for apples(Cheng and Schupp, 2004; Sallato, 2017).
@@ 3.1 Tree Segmentation and Clustering
Figure 12: The images (top row) and segmented point cloud (bottom row) of a test tree during weeks 1,3, and 5. The three colors in the point cloud (bottom row) represent three groups; green-green leaves, yellow-yellow leaves, and brown-trunk, branches and some leaves that have turned brown.
@@ 3.2 Yellowness Index and Leaf Nitrogen Concentration
The correlation between leaf nitrogen concentration and yellowness index was determined for each week in 2021 and 2023. The trees were categorized into five groups based on their leaf nitrogen concentrations. In general, tree nitrogen levels were poorly correlated with the yellowness index in both years (r values ranging from -0.28 to -0.43 in 2021 and 0.16 to -0.64 in 2023). However, an extended retention of greenness was observed in trees that had high N content compared to those with lower N levels. Previous studies have shown strong correlations between color indices and leaf nitrogen concentration in various plant species. For example, Lee et al. (2019) reported a correlation coefficient of 0.85–0.91 between SPAD readings and leaf nitrogen concentration in apples, while El-Azazy (2018) found an R² of 0.81 for citrus trees. Haider et al. (2021) observed an even higher $R^{2}$ of 0.91 across multiple crops, though their system required crop-specific calibration.
Additional studies using hyperspectral imaging have also demonstrated higher predictive accuracy. Ye et al. (2020) reported R² values of 0.77–0.78 in estimating leaf nitrogen concentration using both PLS and MLR models, based on raw reflectance and its first derivative, at both leaf and canopy levels. Jang et al. (2023) used hyperspectral imaging combined with SVR, PLSR, and XGBoost, and maintained similar predictive performance even when reducing spectral resolution through binning strategies, with R² values comparable to full-spectrum models. Li et al. (2022) achieved a validation R² of 0.75 using UAV-acquired hyperspectral imagery and a backpropagation neural network, indicating the potential of aerial platforms for canopy-level nitrogen mapping. Similarly, Chen et al. (2020) reached an R² of 0.843 by combining canopy hyperspectral reflectance with ensemble learning techniques, using extensive preprocessing including SNV-FD and feature selection methods . Wen et al. (2018) also reported an $R^{2}$ of 0.72 using support vector machines on SG-smoothed first derivative spectral data of apple leaves. More recently, Chen et al. (2024) developed a deep learning framework that incorporated phenological data - days after anthosis along with hyperspectral reflectance to estimate leaf nitrogen content, achieving validation R² improvements from 0.69 to 0.79 while eliminating 96% of redundant spectral bands.
In contrast, the method evaluated in this study achieved an $R^{2}$ of 0.72 for estimating the yellowness index using a gradient boost model. Although the correlation between the yellowness index and leaf nitrogen was generally lower than in prior studies, moderate correlations (r = 0.4–0.6) were found during certain weeks, suggesting that specific time windows exist where color-based estimation of nitrogen status may be more effective. The reduced performance in this study may be attributed to factors such as variable natural lighting, occlusion within the canopy, and a relatively narrow distribution of nitrogen concentration values —particularly in the lower N range.
Nonetheless, the system developed here captured relative differences and temporal trends in canopy color across large numbers of trees and under field conditions. Unlike many hyperspectral approaches, which require costly instrumentation and intensive spectral calibration, this system provides a low-cost, scalable, and non-destructive alternative for monitoring nitrogen dynamics. While higher $R^{2}$ values were reported in studies relying on hyperspectral reflectance, the integration of point cloud segmentation and color-based metrics in this study offers unique benefits in field-based decision support contexts. The potential utility of the method can be enhanced by combining the yellowness index with other features such as canopy density, trunk diameter, and phenological data to improve nitrogen estimation in orchards.
@@
The results from this study demonstrated that a machine vision system could be used to identify differences in the yellowing patterns of trees. The outcomes and methodologies of this study offer a means to quantify tree color that could be used to determine tree nitrogen levels, which are crucial factors in guiding fertilizer application to individual trees. However, the system relies on a machine vision system which is susceptible to variations in external lighting conditions. Nevertheless, recent advancements in machine vision sensors have addressed low-light conditions using innovative techniques such as high dynamic range (Zhang et al., 2020) and Adversarial Networks(Chen et al., 2018). In addition, deep learning methods have notably enhanced imaging system resilience to diverse lighting environments(Li et al., 2021). The algorithm employed in this study relies on labeled data for initially training the model to differentiate yellow and green points on trees, necessitating manual inputs. It’s important to note that this algorithm might require further calibration between different years and across tree varieties for optimal performance.